The length of AB in Oo with raduis 3 and m AOB=45
Answer by Boreal(15235)

The central angle is 45 degrees, and that is the arc.
(45/360)=(1/8)
The arc's length is 1/8 of the circumference, and the latter is 2*pi*r=6 pi
Arc's length is (3/4) pi units, which is acceptable, or 2.36 units
The (3/4) comes from 6pi*1/8 of circumference.